# SVD-LLM: Truncation-aware Singular Value Decomposition for Large Language Model Compression

### v2

@@ Abstract
The advancements in Large Language Models (LLMs) have been hindered by their substantial sizes, which necessitate LLM compression methods for practical deployment. Singular Value Decomposition (SVD) offers a promising solution for LLM compression. However, state-of-the-art SVD-based LLM compression methods have two key limitations: truncating smaller singular values may lead to higher compression loss, and the lack of update on the remaining model parameters after SVD truncation. In this work, we propose SVD-LLM, a new SVD-based LLM compression method that addresses the limitations of existing methods. SVD-LLM incorporates a truncation-aware data whitening strategy to ensure a direct mapping between singular values and compression loss. Moreover, SVD-LLM adopts a layer-wise closed-form model parameter update strategy to compensate for accuracy degradation caused by SVD truncation. We evaluate SVD-LLM on a total of $11$ datasets and seven models from three different LLM families at four different scales. Our results demonstrate the superiority of SVD-LLM over state-of-the-arts, especially at high model compression ratios. The source code is available at https://github.com/AIoT-MLSys-Lab/SVD-LLM.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in a wide range of tasks such as natural language understanding and language generation (Zhao et al., 2023; Gozalo-Brizuela & Garrido-Merchán, 2023). Despite such capabilities, the democratization of LLMs is primarily restricted by their substantial resource demands (Wan et al., 2023; Wang et al., 2024). As an example, GPT-3 (Brown et al., 2020) consumes more than 325 GB of GPU memory to load its weights and requires 5×80GB A100 GPUs to conduct inferences due to its high runtime memory consumption (Xiao et al., 2023).
@@
One of the most effective techniques to reduce the resource demands of LLMs is model compression (Zhu et al., 2023). Compression techniques based on quantization (Frantar et al., 2022; Lin et al., 2023; Xiao et al., 2023), parameter pruning (Ma et al., 2023; Frantar & Alistarh, 2023), and knowledge distillation (Gu et al., 2023; Hsieh et al., 2023) specifically designed for LLMs have been intensively studied. Regardless of their successes, these techniques have their own constraints such as hardware dependency and the need of expensive retraining. Compared to those techniques, compression techniques based on low-rank approximation such as Singular Value Decomposition (SVD) are not limited by those constraints. Moreover, the KV cache of LLMs compressed via SVD at runtime can also be reduced.
@@
Despite these advantages, the potential of SVD for LLM compression has not been thoroughly explored. A few SVD-based LLM compression methods such as ASVD (Yuan et al., 2023) and FWSVD (Hsu et al., 2022) have recently been proposed. However, these methods exhibit severe performance degradation when the model compression ratio is high. Such limitation can be attributed to two fundamental issues invovled in their approaches: ❶ Imprecise Data Preprocessing: although the data preprocessing strategy proposed by ASVD reduces the negative impact of activation outliers, it does not establish a direct relationship between SVD approximation and model compression loss. As a consequence, truncating smaller singular values in SVD could lead to significant compression loss. ❷ Lack of Model Parameter Update after SVD Truncation: as the model compression ratio increases, the number of singular values that need to be truncated in SVD increases as well. To compensate for the accuracy degradation caused by truncating a large number of singular values, updating the remaining model parameters to adapt to the new activation is needed. Unfortunately, existing SVD-based LLM compression methods do not take such update into account, and thus fail to compensate for the accuracy degradation under high model compression ratios.
@@
In this paper, we propose a new SVD-based LLM compression method named SVD-LLM that effectively addresses the two fundamental issues of the existing methods. SVD-LLM differs from existing SVD-based LLM compression methods in two key aspects: ❶ Truncation-Aware Data Whitening: Supported by the theoretical proof, SVD-LLM incorporates a truncation-aware data whitening technique that ensures a direct mapping between singular values and model compression loss. In doing so, the proposed truncation-aware data whitening technique is able to identify which singular values should be truncated to incur minimal model compression loss. ❷ Layer-Wise Closed-Form Model Parameter Update: to compensate for accuracy degradation under high compression ratios, SVD-LLM incorporates a layer-wise closed-form model parameter update strategy to progressively update the remaining weights after SVD truncation layer by layer.
@@
We compare SVD-LLM with three SVD-based methods for LLM compression, including vanilla SVD as well as state-of-the-art methods FWSVD and ASVD. To demonstrate the generability of SVD-LLM, we conduct our evaluation on a total of $11$ datasets and seven models from three different LLM families (LLaMA, Vicuna, and Mistral) at four different scales (7B, 13B, 30B, and 65B). We highlight five of our findings: (1) SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all $11$ datasets, three different LLM families, and four different scales. (2) None of the existing SVD-based LLM compression methods report their performance on model compression ratios above 20%. SVD-LLM exhibits significant advantages over vanilla SVD, FWSVD, and ASVD under high compression ratios from 30% to 60%. (3) SVD-LLM also exhibits superiority over state-of-the-art in terms of compression speed. Specifically, when compressing LLaMA-7B under 20% compression ratio on an A100 GPU, ASVD takes about $5.5$ hours whereas SVD-LLM completes the compression process in $15$ minutes. (4) Moreover, SVD-LLM can benefit other LLM compression methods. Our evaluation results show that SVD-LLM is able to further enhance the compression performance of well-recognized quantization (GPTQ (Frantar et al., 2022)) and parameter pruning-based (LLM-Pruner (Ma et al., 2023)) LLM compression methods. (5) Lastly, SVD-LLM brings additional benefit beyond compressing the sizes of LLMs, and is also able to reduce the footprint of KV cache during inference at runtime.
2 Related Works
2.1 Large Language Model Compression
@@
LLMs in general contain billion-scale parameters. Applying conventional model compression methods for LLMs is not feasible as they necessitate retraining. To avoid retraining, post-training methods that do not involve retraining LLM in the compression process have been developed. In general, these methods can be grouped into four categories: unstructured pruning, structured pruning, quantization, and low-rank approximation. Specifically, unstructured pruning methods set the individual weights’ elements to zero without changing its shape. A notable contribution is SparseGPT (Frantar & Alistarh, 2023) which prunes the least important weights’ elements with the inversion of the Hessian matrix. However, the irregular sparsification of unstructured pruning is difficult to achieve the desired speedup or memory saving and can only demonstrate its best efficiency on certain hardware architecture such as NVIDIA Ampere GPU. Unlike unstructured pruning, structured pruning methods directly remove the entire channels or other structured components from LLMs, making them easier to implement on hardware. For example, LLM-Pruner (Ma et al., 2023) utilizes a small amount of data to obtain the weight, parameter, and group importance of the coupled structure for pruning with LoRA to recover the precision. However, due to the great modification of the weight matrix in LLM, it suffers from great accuracy degradation, especially under high compression ratios. Quantization methods, on the other hand, achieve model compression by reducing the precision of weight matrices of an LLM. For example, GPTQ (Frantar et al., 2022) employs layer-wise quantization and updates weights with inverse Hessian information. However, quantization has a drawback of only providing a limited range of compression options, typically ranging from 3 to 8 bits. This limited range could prevent the full utilization of the available memory budget.
@@
2.2 SVD for LLM Compression
Singular Value Decomposition (SVD) is a widely used technique for reducing matrix size by approximating a matrix with two smaller low-ranking matrices (Golub et al., 1987). In the context of LLM compression, only a few SVD-based LLM compression methods have been proposed. Specifically, vanilla SVD only focuses on the compression of the original weight matrix without consider the importance of the parameters, potentially giving a larger compression error. To address this problem, (Hsu et al., 2022) propose FWSVD, which introduces Fisher information to weigh the importance of parameters. However, FWSVD requires complex gradient calculation that demands substantial resources for LLM compression. Another problem of vanilla SVD is the distribution of activation can affect the compression error. To address this issue, (Yuan et al., 2023) propose ASVD, which scales the weight matrix by a diagonal matrix that represents the impact of input channels on weights. However, both FWSVD and ASVD do not establish a direct relationship between singular values and compression errors. As a result, truncating the smallest singular values may not lead to low compression error. Moreover, as the number of truncated singular values increases, the weight matrices in the compressed LLM lose their ability to adjust to new activations. Consequently, these methods incur severe accuracy degradation under high compression ratios.
@@ 3 SVD-LLM
Figure 1 provides an overview of the proposed SVD-LLM. SVD-LLM has two main modules: (1) truncation-aware data whitening, and (2) layer-wise closed-form update. As an overview, before compressing LLM, SVD-LLM collects a small random calibration set following the standard procedure (Frantar & Alistarh, 2023; Yuan et al., 2023; Xiao et al., 2023) to provide data for both truncation-aware data whitening and layer-wise closed-form update. For the data whitening process, SVD-LLM obtains the truncation-aware data whitening matrix by calculating the Cholesky decomposition of the multiplication of the input activation and its transpose. After whitening the input activation, SVD-LLM performs SVD and truncates the weight matrices to compress the LLM. Under high model compression ratios, SVD-LLM performs a layer-wise closed-form update to progressively update the remaining weights layer by layer after compression. In the following, we describe the two main modules truncation-aware data whitening, and layer-wise closed-form update in detail.
@@ 3.1 Truncation-Aware Data Whitening
Motivation: Due to high variance of the input activation, simply applying vanilla SVD for LLM compression leads to severe accuracy degradation (Yuan et al., 2023). To address this issue, ASVD (Yuan et al., 2023) formulates LLM compression as an optimization problem with the following optimization objective:
$O_{1}$ $=\min(||WX-W^{\prime}X||_{F})$ (1)
@@
where $W$ is the original LLM weight matrix, $X$ is the original input activation of $W$, $W^{\prime}$ is the LLM weight matrix after compression, and $L=||WX-W^{\prime}X||_{F}$ is the compression loss in the form of Frobenious loss.
@@
Specifically, ASVD extracts a diagonal matrix $S_{0}$ from $X$ and inserts $S_{0}S_{0}^{-1}$ into the multiplication of $WX$ to normalize the variance of the input activation, i.e, $WX=(WS_{0})(S_{0}^{-1}X)=\tilde{W_{0}}\tilde{X_{0}}$. Subsequently, SVD is performed on the matrix $\tilde{W_{0}}$ instead of $W$ to obtain the decomposed matrices $\tilde{U_{0}}$, $\tilde{\Sigma_{0}}$, and $\tilde{V_{0}}$. Finally, ASVD applies the truncation of the smallest singular values in $\tilde{\Sigma_{0}}$ to obtain the compressed weight matrix $W^{\prime}_{0}=\tilde{U_{0}}\times\text{Trunc.}(\tilde{\Sigma_{0}})\times\tilde{V_{0}}$.
@@
Although reducing the variance of activation improves the performance, ASVD does not establish a direct relationship between singular values and compression loss (a detailed proof is included in Section A.1). To better illustrate this point, we show two concrete examples in Figure 2(a). In the first example ❶ where one singular value is truncated, truncating the smallest singular value 0.1 results in a higher compression loss (loss $=$ 1.1) than truncating 0.9 (loss $=$ 0.7). In the second example ❷ where multiple singular values are truncated, truncating the smallest two singular values 0.9 and 0.1 also leads to a higher loss (loss $=$ 1.9) than truncating 2.4 and 0.1 (loss $=$ 1.7). Therefore, truncating the smallest singular values does not leads to minimal compression loss.
(a) ASVD.
(b) SVD-LLM.
Figure 2: Comparison of data whitening methods between ASVD and SVD-LLM.
@@
Key Design: The key idea of SVD-LLM is to incorporate a truncation-aware data whitening technique that ensures a direct mapping between singular values and compression loss. To achieve this, SVD-LLM change the whitened activation $\tilde{X}=S^{-1}X$ to be orthogonal so that each channel is independent of each other, i.e., $\tilde{X}\tilde{X}^{T}=S^{-1}XX^{T}(S^{-1})^{T}=I$. Though the matrix $S$ is not unique, we found that the Cholesky decomposition (Meyer, 2000) of $XX^{T}$ is a good solution due to its numerical stability. After that we perform SVD on $\tilde{W}=WS$ instead of the original weight $W$ to obtain three decomposed matrices $\tilde{U},\tilde{\Sigma},\tilde{V}$, where $\tilde{U}=[\tilde{u_{1}},\tilde{u_{2}},\tilde{u_{3}},...,\tilde{u_{r}}]$, $\tilde{\Sigma}=\text{diag}(\tilde{\sigma_{1}},\tilde{\sigma_{2}},\tilde{\sigma_{3}},\cdots,\tilde{\sigma_{r}})$, and $\tilde{V}=[\tilde{v_{1}},\tilde{v_{2}},\tilde{v_{3}},...,\tilde{v_{r}}]$. Lastly, the smallest singular values are truncated to compress $\tilde{\Sigma}$ to $\tilde{\Sigma}_{1}$ and obtain the compressed weight $\tilde{W^{\prime}}=\tilde{U^{\prime}}\tilde{\Sigma}_{1}\tilde{V}^{T}$.
@@
Figure 2(b) illustrates the effect of the proposed truncation-aware data whitening technique. In the first example ❶ where one singular value is truncated, the compression loss is equal to the truncated singular value. In the second example ❷, the compression loss of truncating multiple singular values is equal to the square root of the sum of their squares. As such, under the proposed truncation-aware data whitening technique, truncating the smallest singular values leads to minimal compression loss.
@@
In the following, we provide the theoretical proof on why the proposed truncation-aware data whitening technique ensures a direct mapping between singular values and compression loss under the case of one singular value (Theorem 3.2) and multiple singular values (Corollary 3.3).
@@
Using Lemma 3.1, we obtain the compression loss $L_{i}$ when truncating the $i^{th}$ singular value of $\tilde{W}$ to reduce its rank for compression (the compressed weight is denoted by $\tilde{W}^{\prime}$):
$L_{i}$ $=||(\tilde{W}-\tilde{W}^{\prime})\tilde{X}||_{F}$
$=||\tilde{\sigma}_{i}\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}||_{F}=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}\tilde{X}^{T}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}$ (1)
Since both $\tilde{U}=[\tilde{u}_{1},\tilde{u}_{2},\tilde{u}_{3},...,\tilde{u}_{r}]$ and $\tilde{V}=[\tilde{v}_{1},\tilde{v}_{2},\tilde{v}_{3},...,\tilde{v}_{r}]$ are orthogonal matrices, we have:

### v3

@@ Abstract
The advancements in Large Language Models (LLMs) have been hindered by their substantial sizes, which necessitate LLM compression methods for practical deployment. Singular Value Decomposition (SVD) offers a promising solution for LLM compression. However, state-of-the-art SVD-based LLM compression methods have two key limitations: truncating smaller singular values may lead to higher compression loss, and the lack of update on the compressed weight after SVD truncation. In this work, we propose SVD-LLM, a new SVD-based LLM compression method that addresses the limitations of existing methods. SVD-LLM incorporates a truncation-aware data whitening strategy to ensure a direct mapping between singular values and compression loss. Moreover, SVD-LLM adopts a layer-wise closed-form model parameter update strategy to compensate for accuracy degradation under high compression ratios. We evaluate SVD-LLM on a total of $10$ datasets and eight models from three different LLM families at four different scales. Our results demonstrate the superiority of SVD-LLM over state-of-the-arts, especially at high model compression ratios.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in a wide range of tasks such as natural language understanding and language generation [31, 9]. Despite such capabilities, the democratization of LLMs is primarily restricted by their substantial resource demands [25, 26]. One of the most effective techniques to reduce the resource demands of LLMs is model compression [32]. Compression techniques based on quantization [6, 15, 27], parameter pruning [16, 5], and knowledge distillation [10, 11] specifically designed for LLMs have been intensively studied. Regardless of their success, these techniques have their own constraints, such as hardware dependency and the need for expensive retraining. Compared to those techniques, compression techniques based on low-rank approximation, such as Singular Value Decomposition (SVD) are not limited by those constraints.
@@
Despite these advantages, the potential of SVD for LLM compression has not been thoroughly explored. A few SVD-based LLM compression methods such as ASVD [28] and FWSVD [12] have recently been proposed. However, these methods exhibit severe performance degradation when the model compression ratio [^1] The compression ratio refers to the percentage of parameter reduction achieved through compression. is high. Such limitation can be attributed to two fundamental issues involved in their approaches: ❶ Imprecise Data Preprocessing: although the data preprocessing strategy proposed by ASVD reduces the negative impact of activation outliers, it does not establish a direct relationship between singular values and the model compression loss. As a consequence, truncating smaller singular values in SVD could lead to significant compression loss. ❷ Lack of Model Parameter Update after SVD Truncation: as the model compression ratio increases, the number of singular values that need to be truncated in SVD increases as well. To compensate for the accuracy degradation caused by truncating a large number of singular values, it is required to update the remaining parameters in the compressed model. Unfortunately, existing SVD-based LLM compression methods do not take such update into account, and thus fail to compensate for the accuracy degradation under high model compression ratios.
@@
In this paper, we propose a new SVD-based LLM compression method named SVD-LLM that effectively addresses the two fundamental issues of the existing methods. SVD-LLM differs from existing SVD-based LLM compression methods in two key aspects: ❶ Truncation-Aware Data Whitening: Supported by the theoretical proof, SVD-LLM incorporates a truncation-aware data whitening technique that ensures a direct mapping between singular values and model compression loss. In doing so, the proposed truncation-aware data whitening technique is able to identify which singular values should be truncated to incur minimal model compression loss. ❷ Layer-Wise Closed-Form Model Parameter Update: to compensate for accuracy degradation under high compression ratios, SVD-LLM incorporates a layer-wise closed-form model parameter update strategy to progressively update the compressed weights layer by layer.
@@
We compare SVD-LLM with three SVD-based methods for LLM compression, including vanilla SVD as well as state-of-the-art methods FWSVD and ASVD. To demonstrate the generability of SVD-LLM, we conduct our evaluation on a total of $10$ datasets and eight models from three different LLM families (LLaMA, OPT, and Mistral) at four different scales (7B, 13B, 30B, and 65B). We highlight six of our findings: (1) SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all $10$ datasets, three different LLM families, and four different scales and even exhibits significant advantages under high compression ratios from 30% to 60%. (2) SVD-LLM also exhibits superiority over the state-of-the-art in terms of compression speed. Specifically, when compressing LLaMA-7B under 20% compression ratio on an A100 GPU, ASVD takes about $5.5$ hours whereas SVD-LLM completes the compression process in $15$ minutes. (3) The independent performance of either of the two key components of SVD-LLM still consistently surpasses the performance of the current state-of-the-art SVD compression method under different compression ratios. (4) SVD-LLM can benefit other LLM compression methods. Our evaluation results show that SVD-LLM is able to further enhance the compression performance of well-recognized quantization (GPTQ [6]) and parameter pruning-based (LLM-Pruner [16]) LLM compression methods. (5) SVD-LLM can ensure inference speedup on both GPU and CPU. It is able to achieve at most 1.7x speedup on GPU and 1.5x speedup on CPU under the 40% compression ratio. (6) Lastly, SVD-LLM brings additional benefit beyond compressing the sizes of LLMs, and is also able to reduce the footprint of KV cache during inference at runtime.
2 Related Work
@@
Large Language Model Compression: LLMs in general contain billion-scale parameters. Applying conventional model compression methods for LLMs is not feasible as they necessitate retraining. To avoid retraining, post-training methods that do not involve retraining LLMs in the compression process have been developed. In general, these methods can be grouped into four categories: unstructured pruning, structured pruning, quantization, and low-rank approximation. Specifically, unstructured pruning methods set the individual weights’ elements to zero without changing its shape. A notable contribution is SparseGPT [5] which prunes the least important weight elements with the inversion of the Hessian matrix. However, the irregular sparsification of unstructured pruning is difficult to achieve the desired speedup or memory saving and can only demonstrate its best efficiency on certain hardware architecture such as NVIDIA Ampere GPU. Unlike unstructured pruning, structured pruning methods directly remove entire channels or other structured components from LLMs, making them easier to implement on hardware. For example, LLM-Pruner [16] utilizes a small amount of data to obtain the weight, parameter, and group importance of the coupled structure for pruning with LoRA to recover precision. However, due to the great modification of the weight matrix in LLM, it suffers from a great accuracy degradation, especially under high compression ratios. Quantization methods, on the other hand, achieve model compression by reducing the precision of weight matrices of an LLM. For example, GPTQ [6] uses layer-wise quantization and updates the weights with inverse Hessian information. However, quantization has the drawback of only providing a limited range of compression options, typically ranging from 3 to 8 bits. This limited range could prevent full utilization of the available memory budget.
@@
SVD for LLM Compression: Singular Value Decomposition (SVD) is a widely used technique to reduce matrix size by approximating a matrix with two smaller low-ranking matrices [8]. In the context of LLM compression, only a few SVD-based LLM compression methods have been proposed. Specifically, vanilla SVD only focuses on the compression of the original weight matrix without considering the importance of the parameters, potentially giving a larger compression error. To address this problem, [12] propose FWSVD, which introduces Fisher information to weigh the importance of parameters. However, FWSVD requires a complex gradient calculation that demands substantial resources for LLM compression. Another problem of vanilla SVD is the distribution of activation can affect the compression error. To address this issue, [28] propose ASVD, which scales the weight matrix by a diagonal matrix that represents the impact of input channels on the weights. However, both FWSVD and ASVD do not establish a direct relationship between singular values and compression loss. As a result, truncating the smaller singular values may lead to higher compression loss. Moreover, as the compression ratio increases, it is necessary to update the compressed weight due to truncating a great number of singular values. However, existing methods have no design for this update and thus incur severe accuracy degradation under high compression ratios.
@@ 3 SVD-LLM
Figure 1 provides an overview of SVD-LLM. At a high level, SVD-LLM is a SVD-based post-training LLM compression method. Specifically, following the standard procedure of post-training LLM compression methods [5, 28, 27], SVD-LLM uses a random set of sentences as calibration data to generate activation for truncation-aware data whitening and layer-wise closed-form update for model compression. SVD-LLM whitens the activation through Cholesky decomposition, and performs SVD to truncate the weight matrices to compress the LLM. Under high model compression ratios, SVD-LLM performs a layer-wise closed-form update to progressively update the remaining weights layer by layer after compression. In the following, we describe both truncation-aware data whitening and layer-wise closed-form update in detail. The pseudocode is provided in Section A.2.
@@ 3.1 Truncation-Aware Data Whitening
Motivation: Due to high variance of the input activation, simply applying vanilla SVD for LLM compression leads to severe accuracy degradation [28]. To address this issue, ASVD [28] formulates LLM compression as an optimization problem with the following optimization objective:
$O=\min(||WX-W^{\prime}X||_{F})$ (1)
@@
where $W$ is the weight matrix of the original LLM, $X$ is the activation of $W$ given an input, $W^{\prime}$ is the compressed weight matrix, and $L=||WX-W^{\prime}X||_{F}$ is the compression loss in the form of Frobenius loss.
@@
Specifically, ASVD extracts a diagonal matrix $S_{0}$ from $X$ where each element in the diagonal is the absolute mean value of each channel. It then uses $S_{0}$ to normalize $X$ and converts $WX$ into $(WS_{0})(S_{0}^{-1}X)$. Subsequently, SVD is performed on $WS_{0}$ to obtain the decomposed matrices $U_{0}$, $\Sigma_{0}$, and $V_{0}$. Lastly, ASVD truncates the smallest singular values in $\Sigma_{0}$ to obtain the compressed weight matrix $W^{\prime}_{0}=U_{0}\times\text{Trunc.}(\Sigma_{0})\times V_{0}\times S_{0}^{-1}$.
@@
Although normalizing the activation improves the performance, ASVD does not establish a direct relationship between singular values and compression loss (a detailed proof is included in Section A.1). To better illustrate this point, we show two concrete examples in LABEL:fig:asvd_whitening. In the first example ❶ where only one singular value is truncated, truncating the smallest singular value 0.1 results in a higher compression loss (loss $=$ 1.1) than truncating the second smallest singular value 0.9 (loss $=$ 0.7). In the second example ❷ where multiple singular values are truncated, truncating the smallest two singular values 0.9 and 0.1 also leads to a higher loss (loss $=$ 1.9) than truncating 2.4 and 0.1 (loss $=$ 1.7). Hence, truncating the smallest singular values does not lead to minimal loss.
@@
Key Design: The key idea of SVD-LLM is to incorporate a truncation-aware data whitening technique that ensures a direct mapping between singular values and compression loss. To achieve this, SVD-LLM enforces the whitened activation $S^{-1}X$ to be orthogonal such that each channel is independent of each other, i.e., $(S^{-1}X)(S^{-1}X)^{T}=S^{-1}XX^{T}(S^{-1})^{T}=I$, where $S$ is derived through Cholesky decomposition [19]. Then we perform SVD on $WS$ to obtain the decomposed matrices $U,\Sigma,V$, where $U=[u_{1},u_{2},u_{3},...,u_{r}]$, $\Sigma=\text{diag}(\sigma_{1},\sigma_{2},\sigma_{3},\cdots,\sigma_{r})$, and $V=[v_{1},v_{2},v_{3},...,v_{r}]$. Lastly, the smallest singular values in $\Sigma$ are truncated to obtain the compressed weight matrix $W^{\prime}=U\times\text{Trunc.}(\Sigma)\times V^{T}\times S^{-1}$.
@@
LABEL:fig:svdllm_whitening illustrates the effect of the proposed truncation-aware data whitening technique. In the first example ❶ where only one singular value is truncated, the compression loss is equal to the truncated singular value. In the second example ❷, the compression loss of truncating multiple singular values is equal to the square root of the sum of their squares. As such, under the proposed truncation-aware data whitening technique, truncating the smallest singular values leads to minimal compression loss.
@@
Below, we provide a theoretical proof on why the proposed truncation-aware data whitening technique ensures a direct mapping between singular values and compression loss in the case of one singular value (Theorem 3.2) and multiple singular values (Corollary 3.3).
@@
Using Lemma 3.1, we obtain the compression loss $L_{i}$ when truncating the $i^{th}$ singular value of $S^{-1}X$ to reduce its rank for compression:
$L_{i}=||(W-W^{\prime})X||_{F}=||\sigma_{i}u_{i}v_{i}^{T}S^{-1}X||_{F}=\sigma_{i}~{}trace(u_{i}v_{i}^{T}S^{-1}XX^{T}(S^{-1})^{T}v_{i}u_{i}^{T})^{\frac{1}{2}}$ (3)
Since both $U=[u_{1},u_{2},u_{3},...,u_{r}]$ and $V=[v_{1},v_{2},v_{3},...,v_{r}]$ are orthogonal matrices, we have:
$v_{i}^{T}v_{i}=u_{i}^{T}u_{i}=I;v_{i}^{T}v_{j}=u_{i}^{T}u_{j}=0,\forall i\neq j;trace(v_{i}v_{i}^{T})=trace(u_{i}u_{i}^{T})=1$ (4)
Theorem 3.2.
If $S$ is the Cholesky decomposition of $XX^{T}$, the compression loss $L_{i}$ equals to $\sigma_{i}$.
Proof.
Since the whitening matrix $S$ is the Cholesky decomposition of $XX^{T}$, we have $SS^{T}=XX^{T}$. We can further infer Equation 3 to obtain:
$L_{i}=||\sigma_{i}u_{i}v_{i}^{T}S^{-1}X||_{F}=\sigma_{i}~{}trace(u_{i}v_{i}^{T}S^{-1}XX^{T}(S^{-1})^{T}v_{i}u_{i}^{T})^{\frac{1}{2}}=\sigma_{i}~{}trace(u_{i}v_{i}^{T}v_{i}u_{i}^{T})^{\frac{1}{2}}=\sigma_{i}\$ (5)
Therefore, $L_{i}$ of truncating $\sigma_{i}$ equals to the singular value $\sigma_{i}$ itself. ∎
Corollary 3.3.
If $S$ is the Cholesky decomposition of $XX^{T}$, truncating the smallest singular values leads to the lowest loss $L$ compared to truncating others.
Proof.
If we truncate $\sigma_{m+1},\sigma_{m+2},\sigma_{m+3},...,\sigma_{r}$ in $\Sigma$ for compression, the square of the loss $L$ is:
$\begin{split}L^{2}=&\left|\left|\sum_{i=m+1}^{r}\sigma_{i}u_{i}v_{i}^{T}S^{-1}X\right|\right|^{2}_{F}=\sum_{j=m+1}^{r}\sum_{i=m+1}^{r}\sigma_{i}\sigma_{j}~{}trace(u_{i}v_{i}^{T}S^{-1}XX^{T}(S^{-1})^{T}v_{j}u_{j}^{T})\\ =&\sum_{i=m+1}^{r}\sigma_{i}^{2}~{}trace(u_{i}v_{i}^{T}S^{-1}XX^{T}(S^{-1})^{T}v_{i}u_{i}^{T})=\sum_{i=m+1}^{r}(L_{i})^{2}=\sum_{i=1}^{k}(\sigma_{i})^{2}\end{split}$ (6)
The squared loss $L^{2}$ is equal to the sum of the squared singular values. Therefore, truncating the smallest singular values $\sigma_{i}$ achieves the lowest compression loss. ∎
3.2 Layer-Wise Closed-Form Update
Motivation: Given the same calibration data as input, the compressed weight matrix $W^{\prime}$ generates a new activation $X^{\prime}$ that is different from $X$ generated by the original weight matrix $W$. As the compression ratio increases, $W$ needs to truncate a larger number of singular values to obtain $W^{\prime}$, thus $X^{\prime}$ deviates further from $X$. Therefore, it becomes necessary to design a strategy for updating $W^{\prime}$ to minimize $||WX^{\prime}-W^{\prime}X^{\prime}||_{F}$. However, existing SVD-based LLM compression methods have no design of parameter update after compression, leading to less competitive performance at high compression ratios.
Figure 3: Layer-Wise Closed-Form Update.